### v1

# Splines-Based Feature Importance in Kolmogorov-Arnold Networks: A Framework for Supervised Tabular Data Dimensionality Reduction

Abstract
High-dimensional datasets require effective feature selection to improve predictive performance, interpretability, and robustness. We propose and evaluate feature selection methods for tabular datasets based on Kolmogorov-Arnold networks (KANs), which parameterize feature transformations through splines, enabling direct access to interpretable importance measures. We introduce four KAN-based selectors (KAN-L1, KAN-L2, KAN-SI, KAN-KO) and compare them against classical baselines (LASSO, Random Forest, Mutual Information, SVM-RFE) across multiple classification and regression tabular dataset benchmarks. Average (over three retention levels: 20%, 40%, and 60%) F1 scores and $R^{2}$ score results reveal that KAN-based selectors, particularly KAN-L2, KAN-L1, KAN-SI, and KAN-KO, are competitive with and sometimes superior to classical baselines in structured and synthetic datasets. However, KAN-L1 is often too aggressive in regression, removing useful features, while KAN-L2 underperforms in classification, where simple coefficient shrinkage misses complex feature interactions. KAN-L2 and KAN-SI provide robust performance on noisy regression datasets and heterogeneous datasets, aligning closely with ensemble predictors. In classification tasks, KAN selectors such as KAN-L1, KAN-KO, and KAN-SI sometimes surpass the other selectors by eliminating redundancy, particularly in high-dimensional multi-class data. Overall, our findings demonstrate that KAN-based feature selection provides a powerful and interpretable alternative to traditional methods, capable of uncovering nonlinear and multivariate feature relevance beyond sparsity or impurity-based measures.
Keywords: Supervised learning, Tabular data, Feature selection, Kolmogorov-Arnold Networks
1 Introduction
High-dimensional datasets pose challenges to both interpretability and computational efficiency in machine learning. Feature selection addresses these issues by identifying and retaining the most informative features. Traditional supervised approaches to tabular data feature selection fall into three broad categories: filters, wrappers, and embedded methods. Filter methods include statistical criteria such as Pearson correlation and mutual information (MI) (Peng et al., 2005), which are fast and model-agnostic but ignore feature interactions and often select redundant features. Embedded methods integrate selection into the learning algorithm itself, such as Least Absolute Shrinkage and Selection Operator(LASSO) (Tibshirani, 1996a), spline-penalized regression (Marx and Eilers, 1996), tree-based models like Random Forests (Breiman, 2001), and gradient-boosted methods such as XGBoost (Chen and Guestrin, 2016), all of which provide model-aware importance scores but may sacrifice interpretability in complex non-linear settings. Wrapper methods, such as Recursive Feature Elimination (RFE) for Support Vector Machines (SVMs) (Guyon et al., 2002), directly search for subsets of features by iteratively training and evaluating models; these often yield strong predictive performance but are computationally expensive and prone to overfitting. Recent heuristic ensemble strategies such as the Minimum Union (Min) method (Radovic et al., 2017) have also been proposed to combine outputs of multiple selectors, improving stability but adding complexity. Finally, unsupervised projection techniques like Principal Component Analysis (PCA) (Jolliffe, 2002) reduce dimensionality by constructing orthogonal linear combinations of the original variables, which improves computational efficiency but discards direct interpretability and neglects target information. Overall, while traditional methods offer a rich toolbox for feature selection, they involve trade-offs among computational cost, predictive performance, stability, and interpretability. Filter methods are fast but myopic, wrappers are accurate but expensive, and embedded methods are model-aware but often opaque. These limitations motivate the exploration of new architectures, such as Kolmogorov-Arnold Networks (KANs), which offer both model-awareness and direct interpretability by parameterizing each feature transformation as a trainable spline.
Kolmogorov-Arnold Networks (KANs) are a recently proposed neural architecture (Liu et al., 2024b) inspired by the Kolmogorov-Arnold representation theorem, which states that any continuous multivariate function can be expressed as a finite superposition of univariate functions and addition. KANs operationalize this by replacing traditional linear weights with trainable spline functions. Each input feature is mapped through edge-specific splines, aggregated by hidden nodes, and then transformed by learnable outer splines. This “weights-as-functions” paradigm offers both strong approximation power with compact models and intrinsic interpretability: irrelevant features tend to learn nearly flat splines, while informative features produce strong nonlinear mappings. As such, KANs naturally provide mechanisms for assessing feature relevance, making them attractive for supervised feature selection in tabular data. Despite its promise, the use of KANs for feature selection remains largely unexplored. The most explicit attempt is that of Zheng et al. (2025), who proposed a KAN-assisted Whale Optimization Algorithm (KAN-WOA) for high-dimensional medical data, where KAN acted as a surrogate fitness evaluator in evolutionary search. While effective in identifying compact biomarker subsets, this approach inherits the computational cost and sensitivity of meta-heuristic optimization and leverages KAN only as a surrogate, leaving its interpretability underutilized. Other preliminary studies point to similar potential: Wang et al. (2025) used spline visualizations to manually prune redundant features without loss of accuracy, and Liu et al. (2024a) employed a two-layer KAN as a feature extractor for sensor data, outperforming CNN-based alternatives with fewer parameters.
Taken together, these works remain proof-of-concept, underscoring both the promising potential of KAN for supervised feature selection, but also highlighting important gaps: the absence of automated, principled methods for ranking or pruning features directly within KAN, and a lack of large-scale evaluations on diverse tabular datasets. Addressing these gaps is critical to establish KAN as a practical tool for supervised feature selection. consequently, we present the first systematic investigation of Kolmogorov-Arnold Networks (KANs) structures as supervised learning feature selectors for tabular data, providing importance score-free leakage via cross validation. We develop and evaluate four complementary KAN-based importance measures: (i) coefficient-based norms of spline weights (KAN-L1 and KAN-L2), which capture the magnitude of each feature’s learned transformations; (ii) a sensitivity integral approach (KAN-SI), which quantifies the expected gradient of the output with respect to each input; and (iii) a knock-out strategy (KAN-KO), which measures the increase in predictive loss when a feature’s contribution is ablated. Last but not least, we benchmark these four KAN-based feature selectors against six state-of-the-art filter, wrapper, and embedded techniques across eight real-world datasets, assessing predictive performance, selection stability, interpretability, and robustness to irrelevant variables.
The paper is organized as follows. Section 2 surveys related work on feature selection methods. Section 3 provides background and introduces Kolmogorov-Arnold Networks (KANs). Section 4 presents the methodology: coefficient-based feature importance (KAN-L1, KAN-L2), KAN-Knockout (KAN-KO), and KAN-Sensitivity Integral (KAN-SI), and it specifies the assessment pipeline from feature selection to prediction (predictors, evaluation metrics, and leakage-safe cross-validation) together with the baseline methods. Section 5 describes the dataset, experimental results and analysis along with a discussion of different results.
2 Related Work
In supervised learning with tabular data, feature selection is a critical step for improving predictive performance, enhancing interpretability, and reducing computational cost. High-dimensional datasets, common in domains such as finance, bioinformatics, and environmental monitoring often contain irrelevant or redundant variables that can degrade model accuracy and increase overfitting risk. Over time, a variety of supervised feature selection techniques have been developed.
One of the most widely used approaches is the Least Absolute Shrinkage and Selection Operator (LASSO) (Tibshirani, 1996b), which adds an $L^{1}$ penalty to the regression objective to shrink unimportant coefficients to zero. LASSO produces sparse, interpretable models and is computationally efficient, making it attractive for high-dimensional tabular problems. However, when predictors are highly correlated, it tends to retain only one from the group, potentially discarding other relevant variables (Zou and Hastie, 2005). To address LASSO’s inability to model complex nonlinear relationships while enabling flexible yet interpretable modeling of nonlinear effects, Generalized Additive Models (GAMs) (Wood and Augustin, 2002) were proposed for variable selection by fitting penalized splines to the covariates. However, this process is computationally demanding especially when extended to very high-dimensional feature spaces. Tree-based ensemble methods, particularly Random Forests (Breiman, 2001), provide feature importance scores based on mean decrease in impurity or accuracy and can naturally capture feature interactions. While more robust to multicolinearity than linear models, their importance measures can be biased toward variables with more categories or greater variance, and the rankings may not always transfer well to other model types (Strobl et al., 2007). Wrapper approaches such as Recursive Feature Elimination (RFE) (Guyon et al., 2002) attempt to overcome such biases by iteratively training a model, removing the least important features, and repeating until an optimal subset remains. This approach can yield strong predictive accuracy and adapt to various learners, but its repeated training cycles make it computationally expensive for large tabular datasets or complex models. Information-theoretic methods like Mutual Information (MI) ranking (Peng et al., 2005) take a model-agnostic route, measuring the dependency between each feature and the target variable. MI can detect nonlinear dependencies missed by purely linear methods and avoids repeated model training. However, because it evaluates features individually, it ignores redundancy, often selecting correlated variables that add little marginal information.
Introduced by Chen and Guestrin (2016), Extreme Gradient Boosting (XGBoost) rapidly became one of the most widely used tree-ensemble methods, valued for its computational efficiency and built-in feature importance metrics. Tree-based models like XGBoost naturally perform embedded feature selection by evaluating candidate splits during training, and they produce importance scores (e.g., information gain or split frequency) that can be used to rank and filter features. Over the past decade, this capability has been extensively leveraged. For instance, Kasongo and Sun (2020) applied XGBoost to a network intrusion dataset, reducing the dimensionality from 42 to 19 features based on importance scores and achieving substantial gains in classification accuracy. Building on this, Lundberg and Lee (2017) introduced SHAP (SHapley Additive exPlanations), a unified and theoretically grounded framework for quantifying each feature’s contribution to a model’s predictions. SHAP-based feature selections typically rank variables by the magnitude of their contribution to the prediction, providing greater interpretability than native importance scores. These approaches have been adopted in diverse applications for example, in medical settings to identify clinically relevant predictors while also explaining model decisions. Empirical comparisons, however, suggest trade-offs: Wang et al. (2024) found that simple importance-based selection using XGBoost’s native scores often outperformed SHAP-based selection in terms of both AUPRC and efficiency on credit fraud data, given that SHAP computation incurs additional overhead.
Minimum Union (Min) Method (Seijo-Pardo et al., 2019) aggregates multiple feature ranking results to improve stability in high-dimensional tabular data. The Min method computes each feature’s best (lowest) rank across an ensemble of selectors, yielding a combined ranking that favors features identified as important by any selector (Seijo-Pardo et al., 2016). However, the evaluation of Min has largely been confined to such high-dimensional biological data, and it has not been extensively benchmarked on diverse datasets or against a wide range of modern feature selection methods. Thus, while Min can improve selection stability and performance in some scenarios, its general efficacy across other domains and in combination with different learning algorithms remains an open question in the literature.
3 Background
3.1 Kolmogorov-Arnold Networks
Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024b; Akazan et al., 2025) are inspired by Kolmogorov’s superposition theorem, which states that any multivariate continuous function $f(x_{1},\dots,x_{n})$ can be represented as a finite composition of continuous univariate functions. In the KAN formulation, a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is modeled as
$f(\mathbf{x})=\sum_{j=1}^{2n+1}\phi_{j}\!\left(\sum_{i=1}^{n}\psi_{ij}(x_{i})\right),$ (1)
where $\psi_{ij}:[0,1]\to\mathbb{R}$ are univariate feature transformations (implemented via spline expansions), and $\phi_{j}:\mathbb{R}\to\mathbb{R}$ are continuous nonlinearities. Each input feature $x_{i}$ is expanded in a spline basis as
$\psi_{ij}(x_{i})\;=\;\sum_{k=1}^{K}w_{ijk}\,B_{k}(x_{i}),$ (2)
where $\{B_{k}\}_{k=1}^{K}$ are fixed basis functions (e.g., B-splines), and $w_{ijk}$ are trainable coefficients. These parameters are learned end-to-end by backpropagation.
KAN layer (matrix form).
Let $x\in\mathbb{R}^{d}$ be an input vector and $m$ the number of outputs. In this case, a KAN layer computes
$y\;=\;W_{\text{base}}\,\phi(x)\;+\;\sum_{j=1}^{d}W_{\text{spline}}^{(j)}\,b_{j}(x_{j})\;\;\in\;\mathbb{R}^{m},$ (3)
where
  • $W_{\text{base}}\in\mathbb{R}^{m\times d}$ is the base weight matrix,
  • $\phi:\mathbb{R}\to\mathbb{R}$ is an activation applied component-wise to $x$,
  • $b_{j}(x_{j})\in\mathbb{R}^{K}$ is the B-spline basis expansion of $x_{j}$,
  • $W_{\text{spline}}^{(j)}\in\mathbb{R}^{m\times K}$ are the spline weights for feature $j$.
Stacking all spline bases into one long vector
$b(x)\;=\;\big[\,b_{1}(x_{1})^{\top}\;\;b_{2}(x_{2})^{\top}\;\;\cdots\;\;b_{d}(x_{d})^{\top}\big]^{\top}\in\mathbb{R}^{dK}$ (4)
and flattening spline weights into a single matrix
$W_{\text{spline}}\;=\;\big[\,W_{\text{spline}}^{(1)}\;\;W_{\text{spline}}^{(2)}\;\;\cdots\;\;W_{\text{spline}}^{(d)}\,\big]\;\in\;\mathbb{R}^{m\times dK},$ (5)
the layer reduces to
$y\;=\;W_{\text{base}}\,\phi(x)\;+\;W_{\text{spline}}\,b(x).$ (6)
Batch form.
For $X\in\mathbb{R}^{n\times d}$ (rows are samples), define $\Phi(X)\in\mathbb{R}^{n\times d}$ with $(\Phi(X))_{ij}=\phi(X_{ij})$. For the spline part, each feature column $X_{j}$ is expanded into its $K$-dimensional basis $b_{j}(X_{j})\in\mathbb{R}^{n\times K}$. Concatenating these expansions across all features gives
$B(X)\;=\;\big[\,b_{1}(X_{1})\;\;b_{2}(X_{2})\;\;\cdots\;\;b_{d}(X_{d})\,\big]\;\in\;\mathbb{R}^{n\times dK}.$ (7)
Then the outputs $Y\in\mathbb{R}^{n\times m}$ are
$Y\;=\;\Phi(X)\,W_{\text{base}}^{\top}\;+\;B(X)\,W_{\text{spline}}^{\top}.$ (8)
4 Methodology
We propose a family of feature importance measures derived directly from Kolmogorov-Arnold Networks (KANs). Specifically, we consider four complementary approaches: (i) coefficient-based importance using $L^{1}$ and $L^{2}$ norms of spline weights (KAN-L1, KAN-L2), (ii) knock-out importance based on loss differentials (KAN-KO), and (iii) sensitivity-based importance computed from input gradients (KAN-SI).
4.1 Coefficient-Based Feature Importance (KAN-L1 and KAN-L2)
Recall from Eq.( 3) that each feature $x_{j}$ in a KAN layer is associated with a spline weight block
$W_{\text{spline}}^{(j)}\;\in\;\mathbb{R}^{m\times K},$ (9)
where $m$ is the number of outputs and $K$ the number of spline basis functions. This block collects all coefficients linking feature $j$ to every output unit.
The coefficient-based feature importance is then defined by the norm of $W_{\text{spline}}^{(j)}$:
$I_{L_{1}}(x_{j})\;=\;\|W_{\text{spline}}^{(j)}\|_{1}=\sum_{p=1}^{m}\sum_{k=1}^{K}|w_{pjk}|,\qquad I_{L_{2}}(x_{j})\;=\;\|W_{\text{spline}}^{(j)}\|_{2}=\left(\sum_{p=1}^{m}\sum_{k=1}^{K}w_{pjk}^{2}\right)^{1/2}.$ (10)
Finally, o enable comparability across features, these scores are normalized:
$\tilde{I}_{L_{1}}(x_{j})\;=\;\frac{I_{L_{1}}(x_{j})}{\sum_{i=1}^{d}I_{L_{1}}(x_{i})},\qquad\tilde{I}_{L_{2}}(x_{j})\;=\;\frac{I_{L_{2}}(x_{j})}{\sum_{i=1}^{d}I_{L_{2}}(x_{i})}.$ (11)
Extension to base and spline coefficients.
While the definitions above use only the spline weight matrices $W_{i}$, the same construction can be applied to the base weight matrix $W_{\text{base}}\in\mathbb{R}^{m\times d}$. In this case, the $i$-th column of $W_{\text{base}}$ represents the linear contribution of feature $x_{i}$, and its $L^{1}$ or $L^{2}$ norm can be taken as an additional importance score. Combining the norms of both $W_{\text{base}}$ and $W_{\text{spline}}$ yields a measure of the total contribution of each feature, whereas restricting to $W_{\text{spline}}$ isolates the nonlinear component of feature importance.
4.2 KAN-Knockout (KAN-KO)
Let the trained KAN layer (see Eq.( 3)) be parameterized by $W:=(W_{\text{base}},W_{\text{spline}})$. For a given feature index $j\in\{1,\dots,d\}$, we define the knockout operator $\mathcal{K}_{j}$ acting on the first KAN layer by
$\mathcal{K}_{j}(W)\;:=\;\big(\;\widetilde{W}_{\text{base}},\;\widetilde{W}_{\text{spline}}\;\big),\quad\text{where}\quad\widetilde{W}_{\text{base}}(:,j)=0,\;\;\;\widetilde{W}_{\text{spline}}^{(j)}=0.$ (12)
Equivalently, at the level of the layer output
$l^{(-j)}(X)=l(X)-\big[W_{\text{base}}(:,j)\,\phi(X_{j})+W_{\text{spline}}^{(j)}\,b_{j}(X_{j})\big].$ (13)
To quantify the contribution of each input feature, we evaluate how much the task loss increases when that feature is removed from the model. Let $\ell(\hat{Y},Y)$ denote the task loss and $P$ the data distribution over $(X,Y)$. We write $f_{W}$ for the full KAN model, where $W$ collects the base and spline weights (see  3.1). The expected or population risk of the model is
$L(W)\;=\;\mathbb{E}_{(X,Y)\sim P}\!\big[\,\ell(f_{W}(X),Y)\,\big].$ (14)
For a given feature index $j$, we define a knock-out operator $\mathcal{K}_{j}(W)$ that sets to zero the $j$th column of $W_{\text{base}}$ and the $j$th spline block of $W_{\text{spline}}$. The corresponding risk when feature $j$ is removed is
$L_{j}(W)\;=\;\mathbb{E}_{(X,Y)\sim\mathcal{D}}\!\big[\,\ell(f_{\mathcal{K}_{j}(W)}(X),Y)\,\big].$ (15)
The KAN-KO importance score for feature $j$ is the nonnegative increase in risk,
$\Delta_{j}\;=\;\max\!\big\{0,\;L_{j}(W)-L(W)\big\}.$ (16)
To make scores comparable across features, we normalize them as
$\tilde{I_{j}}\;=\;\frac{\Delta_{j}}{\sum_{k=1}^{d}\Delta_{k}+\delta},$ (17)
where $\delta>0$ is a small constant to avoid division by zero.
4.3 KAN- Sensitivity Integral (KAN-SI) Feature Importance
For an input feature $x_{i}$, the instantaneous (local) sensitivity of $f$ is the magnitude of its partial derivative,
$S_{i}(\mathbf{x})\;=\;\left|\frac{\partial f(\mathbf{x})}{\partial x_{i}}\right|.$ (18)
The KAN-SI importance is the (data) expectation of $S_{i}$:
$I_{i}=\mathbb{E}_{\mathbf{x}\sim P}[S_{i}(\mathbf{x})],\hskip 14.22636pt\widehat{I_{i}}=\dfrac{1}{N}\sum_{n=1}^{N}\left|\frac{\partial f(\mathbf{x^{n}})}{\partial x_{i}}\right|$ (19)
where $\widehat{I}_{i}$ is the empirical estimator over a held-out set $\{\mathbf{x}^{n}\}_{n-1}^{N}$ (validation or out-of-fold) to avoid optimistic bias. From (1), let the inner pre-activations be
$t_{j}(\mathbf{x})=\sum_{i=1}^{n}\psi_{ij}(x_{i}).$
Then by the chain rule,
$\frac{\partial f(\mathbf{x})}{\partial x_{i}}\;=\;\sum_{j=1}^{2n+1}\phi^{\prime}_{j}\!\big(t_{j}(\mathbf{x})\big)\,\frac{\partial\psi_{ij}(x_{i})}{\partial x_{i}}.$ (20)
Because $\psi_{ij}$ is a spline expansion,
$\frac{\partial\psi_{ij}(x_{i})}{\partial x_{i}}=\sum_{k=1}^{K}w_{ijk}B^{\prime}_{k}(x_{i}),$ (21)
and for B-splines of degree $p$, only the basis derivatives $\big(B^{\prime}_{k}(x_{i})\big)$ of at least degree $p+1$ are nonzero at a given $x_{i}$ (local support), making (20) efficient to evaluate.
Since sensitivities depend on the measurement of $x_{i}$, we report a normalized score.
$\tilde{I}_{i}=s_{i}\widehat{I}_{i},\qquad s_{i}\in\big\{\mathrm{Std}(x_{i}),\mathrm{IQR}(x_{i})\big\}$ (22)
where $\mathrm{Std}$ (standart deviation) or $\mathrm{IQR}$ (interquartile range) is calculated on the same split used for (19). For one - hot encoded categoricals, compute $\tilde{I}$ per dummy and sum back to the original category.
4.4 Assessment Procedure (From Feature Selection to Prediction)
Let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\begin{cases}\{1,\dots,C\},&\text{classification},\\ \mathbb{R},&\text{regression}.\end{cases}$.
Let $\mathcal{S}=\{s_{1},\dots,s_{L=9}\}$ denote the feature selectors. Each $s\in\mathcal{S}$ fitted on training data returns a nonnegative importance vector $a_{s}\in\mathbb{R}_{+}^{d}$, normalized so $\sum_{j=1}^{d}a_{s,j}=1$. For a retention ratio $k\in\{20\%,40\%,60\%\}$, set $n_{k}=\max\!\left\{1,\,\left\lceil\frac{k\,\cdot\,d}{100}\right\rceil\right\},$ representing the number equivalent of k% of the features.
Let $J_{s,k}\subset\{1,\dots,d\}$ be the indices of the $n_{k}$ largest entries of $a_{s}$ (ties arbitrary), and define the projection $\Pi_{J}(x)\in\mathbb{R}^{|J|}$ that restrict $x$ to the matrix defined by the coordinates $J$.
4.4.1 Predictors and evaluation metrics.
For each task $t\in\{\text{classification},\text{regression}\}$, we consider a family of predictors $\mathcal{P}_{t}=\{P_{t,1},\dots,P_{t,M=4}\}$, including Logistic Regression (LogReg) / Ridge, Random Forests (RF), Gradient Boosted Trees (GBT), and XGBoost (XGB). Predictive performance is quantified by
$S_{t}(y,\hat{y})=\begin{cases}\text{Macro-}F_{1}(y,\hat{y}),&t=\text{classification},\\ R^{2}(y,\hat{y}),&t=\text{regression},\end{cases}$ (23)
where the per-class $F_{1}$ score is defined one-vs-rest and averaged across $C$ classes:
$F1_{c}=\frac{2TP_{c}}{2TP_{c}+FP_{c}+FN_{c}},\qquad\text{Macro-}F_{1}=\frac{1}{C}\sum_{c=1}^{C}F1_{c}.$ (24)
4.4.2 Leakage-safe cross-validation.
We use $F$-fold cross-validation with splits $(T_{f},V_{f})$, $f=1,\dots,F$. For each selector $s$, retention level $k$, predictor $P_{t,m}$, and fold $f$:
  1. 1.
    fit preprocessing and selector $s$ on $\{(x_{i},y_{i})\}_{i\in T_{f}}$ to obtain $J_{s,k}^{(f)}$;
  2. 2.
    train $P_{t,m}$ on the projected training set $\{(\Pi_{J_{s,k}^{(f)}}(x_{i}),y_{i})\}_{i\in T_{f}}$;
  3. 3.
    predict on the projected validation set $\{\Pi_{J_{s,k}^{(f)}}(x_{i})\}_{i\in V_{f}}$ to obtain $\{\hat{y}_{i}^{(s,k,t,m,f)}\}_{i\in V_{f}}$.
The fold-level score is
$S_{s,k,t,m,f}=S_{t}\!\Big(\{y_{i}\}_{i\in V_{f}},\,\{\hat{y}_{i}^{(s,k,t,m,f)}\}_{i\in V_{f}}\Big),$ (25)
where $S_{t}$ is macro-$F_{1}$ for classification and $R^{2}$ for regression. The cross-validated score is then
$\mathrm{Score}(s,k,t,m)=\frac{1}{F}\sum_{f=1}^{F}S_{s,k,t,m,f}.$ (26)
4.5 Comparison with Baseline Methods
We compare our spline-based feature selectors against Mutual Information (MI) ranking (Peng et al., 2005), Random Forest (RF) (Breiman, 2001), LASSO-based feature selection (Tibshirani, 1996b), and Recursive Feature Elimination (RFE) for Support Vector Machine (SVM) (SVM-RFE)(Guyon et al., 2002).
5 Experiments
5.1 Data Sets
To evaluate the effectiveness of the Kolmogorov-Arnold Networks dimensinality reduction methods, experiments were conducted on a diverse collection of benchmark datasets covering both classification and regression tasks. These datasets includes well-established repositories from scikit-learn and seaborn as well as synthetically generated datasets and are summarized in Table 1.
Table 1: Benchmark and synthetic datasets
Task Dataset Observations Numbers of features References
Classification Wine 178 13 (Forina et al., 1991).
Breast Cancer Wisconsin 569 30 (Street et al., 1993)
Digits 11,797 64 (Pedregosa et al., 2011)
Make classification 500 10 (Pedregosa et al., 2011)
Regression California Housing 20,640 8 (Efron et al., 2004)
Diabetes 442 10 (Efron et al., 2004).
Diamonds 53,940 10 (Wickham, 2008)
Make regression 500 10 (Pedregosa et al., 2011)
5.2 Results
The comparative analysis across datasets reveals that averaging the macro F1 and $R^{2}$ scores across retention levels $k$ reveals consistent yet dataset-specific interactions between selectors, predictors, and the underlying data structure.
Regression datasets
On the California dataset (nonlinear, heterogeneous), Figure. 1, KAN-L2 and KAN-SI preserve tree-ensemble performance: with Random Forest and XGBoost they stay close to the prediction using all Features (minor drops $\sim$0.01-0.03), while Gradient Boosted Tree is slightly more sensitive. In contrast, KAN-L1 is overly aggressive: it depresses Ridge markedly (large $R^{2}$ gap vs. All Features) and also trims XGBoost and Gradient Boosted Tree more than other KAN variants, indicating loss of weak but complementary signals. KAN-KO feature selector is generally competitive, trailing All Features by a small margin and often matching the classical Mutual Info, Random Forest and SVM-RFE subsets.
Figure 1: Average $R^{2}$ Score (averaged across 20/40/60% retained features) per Selectors for Each Regressor
For Diabetes dataset (small-$n$, noisy), Figure. 1), KAN-L1 again under-selects (lowest bars for Gradient Boosted Tree, Random Forest, XGBoost), whereas KAN-L2, KAN-SI yield the most stable KAN performance across predictors and typically track LASSO and Mutual Info. KAN-KO behaves as a conservative KAN subset, usually second-best among KANs. In the Diamond dataset (strong signal, structured, mixed types), Figure. 1, All models are high-$R^{2}$; here KAN-KO, KAN-L2 and KAN-SI are nearly indistinguishable from All Features for Random Forest, XGBoost and Gradient Boosted Tree and even outperform classical features selectors (Mutual Info, Random Forest and SVM-RFE subsets, LASSO and Random Forest). KAN-L1 is the only KAN variant that noticeably drops on some predictors, consistent with over-pruning informative but correlated attributes. For the Synthetic make regression dataset (well-conditioned, Figure. 1, All KAN variants cluster tightly with baselines across predictors (tiny spreads): when the signal is clean and features are already informative, KAN selection neither helps nor hurts much; KAN-L2/KAN-SI remain the safest choices.
Classification datasets
In the high-dimensional Breast Cancer dataset (Figure 2), Gradient Boosted Trees perform best with KAN-L1, LASSO, and RF importance, often matching the full feature baseline. In contrast, KAN-KO and Mutual Information yield poor subsets, while Logistic Regression and tree ensembles (Random Forest, XGBoost) benefit most from sparsity or impurity-based selectors. This indicates that selectors capturing either linear sparsity or tree-based splits are most effective, whereas knock-out and sensitivity KAN variants are less suited in higher dimensions.
Figure 2: Average Macro $F1$ Score (across 20/40/60% retained features) per Selectors for Each Classifiers
In the high-dimensional Digits dataset (Figure 2), the pattern shifts: Gradient Boosted Trees and XGBoost achieve higher F1 scores with LASSO, SVM-RFE, and several KAN variants (KAN-KO, KAN-SI, KAN-L1), sometimes surpassing the full-feature baseline. Here, dimensionality reduction removes redundancy and sharpens predictive power. By contrast, KAN-L2 and Mutual Information underperform, reflecting their limits in capturing complex feature interactions in high dimension datasets. For the synthetic make classification dataset (Figure 2), all KAN-based selectors (KAN-L1/L2, KAN-SI, KAN-KO) consistently outperform classical baselines, confirming their advantage in structured, high-signal settings where ground-truth feature relevance extends beyond sparsity or univariate dependence. Finally, on the Wine dataset (Figure 2), the advantage shifts back toward classical selectors: RF importance, LASSO, Mutual Information, and SVM-RFE provide the strongest subsets, while most KAN-based selectors lag behind. Only KAN-SI remains competitive, suggesting that when observations are few and feature relationships are well-structured, simpler selectors aligned with linear sparsity or impurity-based measures are more effective.
5.3 Discussion
Across classification and regression, performance hinges on data properties (dimensionality, noise, correlation) and predictor bias. For classification, linear models (Logistic Regression) pair best with sparse and coefficient-driven selectors (LASSO, KAN-L2), whereas ensembles (RF, XGB) benefit from multi-feature selectors (RF importance, KAN-L1/KAN-SI/KAN-KO, SVM-RFE). In high-dimensional sets (Digits), judicious reduction (LASSO, KAN-L1, SVM-RFE) often outperforms the full feature set, revealing harmful redundancy; on synthetic data, all KAN variants recover ground-truth features and rival or beat baselines. For regression, heterogeneous/noisy data (California, Diabetes) favor KAN-L2 and KAN-SI, which track tree ensembles and Ridge with minimal loss, while KAN-L1 is over, aggressive and hurts boosting/linear models. On structured or strong-signal data (Diamond, synthetic), most selectors tie, but KAN-KO/KAN-L2/KAN-SI consistently match or exceed classical baselines.
6 Conclusion
This work presented the first systematic evaluation of Kolmogorov-Arnold Network (KAN) based feature selection across both classification and regression tabular data benchmarks. The results show that KAN selectors, particularly KAN-L2, KAN-SI, and KAN-KO, offer competitive and often superior performance to classical baselines in structured or synthetic settings, while remaining robust on real-world datasets. KAN-L1, though occasionally (for classification datasets) strong, tends to over-prune and degrade predictive power in noisy or correlated contexts especially in regression datasets. Classical selectors such as LASSO and Random Forest remain strong contenders, especially when aligned with linear or ensemble predictors, respectively. Overall, the study highlights that no single selector dominates universally, and that the value of KAN-based feature selection lies in its ability to adaptively capture nonlinear and multivariate dependencies that traditional methods may overlook.